tera traveled 120 miles at a certain speed. tara then drove another 100 miles at a speed that was 10 mph slower. If the total time of taras trip was 4 hours, what was her speed at each part of the trip
Answer by stanbon(75887)

tera traveled 120 miles at a certain speed. tara then drove another 100 miles at a speed that was 10 mph slower. If the total time of taras trip was 4 hours, what was her speed at each part of the trip
------
1st leg DATA:
distance = 120 miles ; speed = x mph ; time = 120/x hrs
----
2nd leg DATA:
distance = 100 miles ; speed = x-10 mph ; time = 100/(x-10) hrs
----
Equation:
120/x + 100/(x-10) = 4 hrs
-----
Modify:
30/x + 25/(x-10) = 1
30(x-10) + 25x = x^2-10x
---
x^2 -10x - 55x + 300 = 0
----
x^2 - 65x +300 = 0
----
(x-60)(x-5) = 0
Positive useful solution:
x = 60 mph
-------------------------------
Cheers,
Stan H.
================